OpenStudy (anonymous):

can some one please help me with these questions ill give a medal for each ques

OpenStudy (chris911):

post the question

OpenStudy (anonymous):

@Chris911

OpenStudy (chris911):

ok so how many degrees are in an rectangle/square??

OpenStudy (anonymous):

90 ?

OpenStudy (chris911):

there is 360 degrees in an square/rectangle

OpenStudy (anonymous):

yea your right i wasn't thinking

OpenStudy (chris911):

hang on @jim_thompson5910 @ParthKohli @Ashleyisakitty I believe it would be 100 or 84 becuase it looks like it has to equal to 360

OpenStudy (anonymous):

which one you think it is

OpenStudy (anonymous):

@Chris911

OpenStudy (chris911):

@zepdrix @ganeshie8 I wanna say its 84 I thinnk

OpenStudy (anonymous):

@Chris911 can you answer one more ques ?

OpenStudy (chris911):

will try if I cannot tag those users I tag earlier^ im a little busy

OpenStudy (anonymous):

@Chris911

OpenStudy (chris911):

hm which one do you think it is?

OpenStudy (chris911):

bunnies this one

OpenStudy (anonymous):

idk

OpenStudy (ipwnbunnies):

Wow, I'm not sure. I tried to get it started but I'm not sure what to do.

OpenStudy (chris911):

I would say 80 since its almost close to 90

OpenStudy (ipwnbunnies):

Wait, I'm not sure which angle B is. >.>

OpenStudy (anonymous):

What is the value of a to the nearest tenth? 7.5 15.2 12.9 17.5 https://wcps6429-millbrookhs-ccl.gradpoint.com/Resource/3750901,7B0,111,0,0,0,0/Assets/testitemimages/geometry_b/circles/mc067-1.jpg

OpenStudy (anonymous):

how abt this one

OpenStudy (ipwnbunnies):

I think this is just a proportion one. \[\frac{15}{30} = \frac{a}{35}\] Solve for a.

OpenStudy (anonymous):

so its 15.2 or 12.9 ... idk

OpenStudy (anonymous):

think i did it wrong

OpenStudy (ipwnbunnies):

You did. To solve for a, multiply both sides by 35. \[\frac{15}{30}*35 = a\]

OpenStudy (chris911):

so it 17.5

OpenStudy (anonymous):

can i ask you another question and ill give you a medal in a new post ?

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Yeah, 17.5. And sure.